OpenStudy (anonymous):

An equation is shown below. 7x+Y=5 Part A: Explain how you will show all of the solutions that satisfy this equation. ` Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation to form a system of equations that is inconsistent.

OpenStudy (anonymous):

@math&ing001 ? lol

OpenStudy (anonymous):

@Its_Aysiaa ? lol

OpenStudy (anonymous):

Sweetheart, I'm HORRIBLE at math sowee :/

OpenStudy (anonymous):

the answer is x=-1/7y+5/7

OpenStudy (anonymous):

@321oky : ok for what part was that?

OpenStudy (anonymous):

you Solve the Equation for x

OpenStudy (anonymous):

I just need the awnsers for all three parts! I suck a math.

OpenStudy (anonymous):

@Jpenrod ?

OpenStudy (anonymous):

part c is You get the equation 7x + y = 5 and multiply it by any number to get a new equation so 7x + y = 5 multiplied by 2 equals 14x + 2y = 10

OpenStudy (anonymous):

well for b I can only think of two different solutions... solving for x and then solving for y. x would equal what 321 got and y would equal y=5-7x

OpenStudy (anonymous):

What about Part a?

OpenStudy (anonymous):

im lost on part A

OpenStudy (anonymous):

Mmm its okay, thank you though!